OpenStudy (anonymous):

Find sin 2theta and cos 2theta if tan theta= square root of two? I think I need to use a double-angle formula but I don't know which one or how to start, please help!

OpenStudy (anonymous):

u can express ..sin2theta and cos2theta in terms of tantheta also ..!

OpenStudy (anonymous):

\[\sin2x=\frac{ 2tanx }{ 1+\tan^2x } ,,\cos2x=\frac{ 1-\tan^2x }{ 1+\tan^2x }\]

OpenStudy (anonymous):

so sin2x=\[\frac{ 2\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

and cos2x=-1/3